# Restless Multi-armed Bandits under Frequency and Window Constraints for Public Service Inspections

### v1

@@ Abstract
Municipal inspections are an important part of maintaining the quality of goods and services. In this paper, we approach the problem of intelligently scheduling service inspections to maximize their impact, using the case of food establishment inspections in Chicago as a case study. The Chicago Department of Public Health (CDPH) inspects thousands of establishments each year, with a substantial fail rate (over 3,000 failed inspection reports in 2023). To balance the objectives of ensuring adherence to guidelines, minimizing disruption to establishments, and minimizing inspection costs, CDPH assigns each establishment an inspection window every year and guarantees that they will be inspected exactly once during that window. These constraints create a challenge for a restless multi-armed bandit (RMAB) approach, for which there are no existing methods. We develop an extension to Whittle index-based systems for RMABs that can guarantee action window constraints and frequencies, and furthermore can be leveraged to optimize action window assignments themselves. Briefly, we combine MDP reformulation and integer programming-based lookahead to maximize the impact of inspections subject to constraints. A neural network-based supervised learning model is developed to model state transitions of real Chicago establishments using public CDPH inspection records, which demonstrates 10% AUC improvements compared with directly predicting establishments’ failures. Our experiments not only show up to 24% (in simulation) or 33% (on real data) reward improvements resulting from our approach but also give insight into the impact of scheduling constraints. [^1] Source code for our algorithms and experiments is available on https://anonymous.4open.science/r/TimeConstrainedBandit-024C/.
@@ 1 Introduction
Restless multi-armed bandits (RMABs) Whittle (1988) are almost a natural fit for this problem. They describe a sequential decision problem where an agent (the city) acts on a large population of independently evolving Markov decision processes (the establishments), which describe each establishment’s propensity to stay in the inspection-passing state. While RMABs are highly suitable for maximizing the inspection-passing probability for each establishment, existing approaches (e.g.,  Yu et al. (2018); Herlihy et al. (2023); Li and Varakantham (2023)) fail to support the constraint structure required for establishment scheduling, where each arm’s pulls must satisfy an ex-post frequency constraint.
@@
  • Empirical evaluation and impact assessment. For the CDPH, we develop a machine-learning approach to estimate the parameters of each establishment based on historical data. We evaluate reward and computation time in both real and synthetic data. In our tested cases, we find that both explicitly modeled constraints and optimized action windows bring substantial advantages compared with ad hoc methods for handling constraints.
  • Machine-learning model predicting transitions of MDP. There is one previous work predicting the food inspection failures directly using machine-learning models Schenk Jr. et al. (2015). For the purpose of building MDPs for our RMAB, we instead train a neural network to predict the state transition probabilities. We found the model not only aligns with the requirements of RMAB, but also has the ability to predicting the inspection failures. It improves the AUC number by around 10% compared with the XGBoost model proposed by Schenk Jr. et al. (2015).
@@ 2 Related Work
In 2015, CDPH leveraged historical food inspection data and trained a supervised learning model to predict the probability that an inspection would uncover a critical violation Schenk Jr. et al. (2015).  Kannan et al. (2019) independently analyzed the impact of prediction-driven scheduling. However, such models only consider one-shot predictions for critical violations and do not include the sequential aspect of scheduling. Fairness is of substantial interest in the provision of municipal services  Singh et al. (2022). We consider fairness outside the scope of this paper, but a potentially interesting direction for future work.
Restless multi-armed bandits (RMABs). RMABs are PSPACE-hard in the worst case, but Whittle (1988) showed that a subclass of them, so-called indexable RMABs, admit an efficient asymptotically optimal solution. Particularly relevant classes of indexable RMABs are those that extend the machine maintenance problem families Glazebrook et al. (2006), and scheduling problems for sensors Sombabu et al. (2020), wireless transmission Hsu (2018), and health interventions Mate et al. (2020). These RMABs are structured so the state of each process declines if it is not acted on, and differ in the details of action effect and what information is observed with or without an action. This work aims to develop techniques to integrate action constraints into RMABs of these types.
RMABs with Constraints. Several RMAB models have included constraints. In a project applying RMABs to assist maternal and child health via phone calls, a “sleeping period” for arms was enforced after they were pulled by the Whittle index heuristic Mate et al. (2022) (See 4.2). It appears to have been enforced in an ad hoc manner, by blocking pulls that would have violated the constraint.  Yu et al. (2018) deployed RMABs in a deadline scheduling setting and integrated the deadline constraints by adding dummy arms. Fairness is another setting where constraints can arise.  Herlihy et al. (2023) introduced the ProbFair policy, ensuring a strictly positive lower bound on the probability of being pulled at each time step while still satisfying the budget constraints.  Li and Varakantham (2023)’s SoftFair balances the trade-off between the goals of having resources uniformly distributed and maximizing cumulative rewards. They guarantee a long-term probability of each arm being pulled whereas ours ensures pulling frequencies for each arm strictly in a period. To the best of our knowledge, we are the first to consider action window and frequency constraints strictly for recurring tasks in a period.
@@ 3 Preliminaries
An RMAB consists of $N$ binary action MDPs (arms). We define the $i$th two-action MDP Puterman (1994) as a tuple $(\mathcal{S}_{i},\mathcal{A},P_{i},R_{i},s^{(0)}_{i},\gamma)$. The discount factor $\gamma$ and action space $\mathcal{A}=\{0,1\}$ are fixed across all MDPs. When the action 1 (resp. 0) is taken on an arm at time $t$, we refer to that arm as active (resp. passive). The rest are arm specific: $\mathcal{S}_{i}$ is the state space, $P_{i}:\mathcal{S}_{i}\times\mathcal{A}\rightarrow\Delta\mathcal{S}_{i}$ is the transition function, $s^{(0)}_{i}$ is the start state, and $R_{i}:\mathcal{S}_{i}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. Because there are only two actions, the transition function $P_{i}$ can be decomposed into an an active transition $P^{(1)}_{i}:\mathcal{S}_{i}\rightarrow\Delta\mathcal{S}_{i}$ and a passive transition $P^{(0)}_{i}:\mathcal{S}_{i}\rightarrow\Delta\mathcal{S}_{i}$.
@@ 3.1 Whittle Indices
General RMABs have an exponentially large state space and a combinatorially large action space. The Whittle index method provides tractability for some classes of RMABs Whittle (1988). It works by computing a “benefit of acting” for each arm, called the Whittle index. The Whittle index heuristic then acts on the $k$ arms with highest Whittle indices.
@@
The goal is to identify the smallest subsidy $m$ such that, for the current state $s_{i,t}$, the long-term reward for the passive and active actions are the same. To formalized this, we first define the $Q$ function for arm $i$ under subsidy $m$:
@@
Definition The Whittle index for state $s_{i,t}$ is the smallest m which makes it equally optimal to take the active and passive actions:
@@
For the Whittle index heuristic to have asymptotic optimality guarantees, each arm must satisfy a technical condition called indexability Whittle (1988). Intuitively, indexability says that, as $m$ increases, the optimal action can only switch to passive and cannot switch back to active. Let $W_{i}^{(m)}$ be the set of states for which $Q_{i}^{(m)}(s_{i,t},a=0)\geq Q_{i}^{(m)}(s_{i,t},a=1)$, i.e., the passive action has an equal or higher return than the active action.
Definition (Indexability). An arm is said to be indexable if $W_{i}^{(m)}$ is non-decreasing in m, i.e., for any $m_{1},m_{2}\in\mathbb{R}$ such that $m_{1}\leq m_{2}$, we have $W_{i}^{(m_{1})}\subseteq W_{i}^{(m_{2})}$. An RMAB is indexable if every arm is indexable.
@@ 4.1 Motivating Inspection RMAB
Motivated by the food establishment setting, we define a model RMAB with action constraints. This RMAB can be viewed as a collapsing bandit Mate et al. (2020) or a resetting bandit Khansa et al. (2021), and both have indexability guarantees. Each establishment has an unobserved binary state that is either 1 (i.e., inspection passing) or 0 (i.e., inspection failing). When we act on the establishment, we assume that it is restored to the passing state and define the reward function to be 1 for each time period the establishment is in the passing state and 0 otherwise. We consider time periods as months—each establishment needs to be inspected once a year and will have a two-month period where this inspection can occur.
As the true states are not directly observable, each arm is a partially observed Markov decision process (POMDP) Araya et al. (2010). We can rewrite the POMDP as a fully observed belief-state MDP, allowing for direct representation as an RMAB.
@@ 4.2 Action Windows and MDP Encoding
We use action windows as an exemplar for the family constraints where the constraint can be directly encoded into the RMAB structure, i.e., a vanilla RMAB with an action window constraint can be rewritten as a vanilla RMAB with a different arm structure. This is in some sense the ideal way to add constraints—we can apply whatever existing state-of-the-art algorithm directly.
@@
As shown in Figure 1, such an encoding increases the number of states in the MDP. The number of states is increased by a factor of $O(LM)$, where $L$ is the number of counter values required to track when the window is active, and $M$ is the total number of actions allowed during the window. For the motivating RMAB, this increase is by a factor of 14. But such encoding is also applicable to more complex situations: An arm has multiple action windows and multiple inspections allowed per window. We provide a precise description of the new MDP in the appendix. All new transitions are deterministic (probability = 1).
To enforce that there are $\eta$ timesteps of no action (sleep) after each action, as in  Mate et al. (2022), requires a factor of $\eta$ more states. We add a counter to the state to record the number of timesteps until the next pull is allowed. When the counter is positive, the effect of an action is the same as the effect of no action.
The Whittle index of arms that are not eligible to be pulled is zero. Arms outside the action window (or during mandatory sleep) have no advantage for the active action over the passive action. Thus, their Whittle index will be zero. In practice, this means the Whittle heuristic will never select these arms, as long as there is some arm with positive action effects. If they are selected anyway, the agent can discard these actions to no ill effect.
@@ 4.3 Frequency Constraints and Lookahead
Maximizing the sum of Whittle indices without additional frequency constraints OR with the constraint that each arm must be pulled exactly once during the lookahead window can be reduced to a weighted $b$-matching.
@@
In practice, it is convenient to solve this lookahead problem as an IP. We can do so with $NL$ binary variables $a_{i,t}$ (where $L$ is the length of the action window) and $T+N$ constraints. Because the polynomial tractability of weighted $b$-matching arises from total unimodularity Schrijver and others (2003), the IP can be solved very quickly via its LP relaxation. However, polynomial tractability is lost when sufficiently complex minimum and maximum number of pulls are added as additional constraints as the problem becomes equivalent to weighted bipartite $b$-matching Chen et al. (2016).
@@ 5 Action Window Optimization
How do we assign windows according to the virtual inspection sequence? Let $a_{i,t}$ be the encoding of the virtual inspection sequence, where $a_{i,t}=1$ if arm $i$ is pulled at time $t$. A naive way is to assign arm $i$ with window $[t,t+W-1]$ if $a_{i,t}=1$. However, such assignments are easily predictable and establishments would be able to prepare effectively. We need to design a way in which establishments get inspected with a probability of $1/W$ on each day in the window.
We can do so with a linear program (LP). We define variables $f_{t,t^{\prime}}$, which indicate the proportion of arms with virtual inspection time $t$ that are put in inspection window $[t^{\prime},t^{\prime}+W-1]$. We denote the number of arms with assigned inspection $t$ that are put in the window starting at $t^{\prime}$ as $g_{t,t^{\prime}}$. Thus, $g_{t,t^{\prime}}=\sum\limits_{i=1}^{N}a_{i,t}f_{t,t^{\prime}}$. For the objective, we use
@@
    The virtual action assignment must occur during the window: $f_{t,t^{\prime}}\neq 0$ if and only if $10\leq t-t^{\prime}\leq W-$ 1.
@@
Using the $f_{t,t^{\prime}}$ output of the LP, we sample windows using any categorical sampling procedure. At this stage, the virtual inspection can be discarded and individual inspections planned as if the windows were given (i.e., not even the system operator knows when the virtual inspections were planned to occur).
@@ 6 Experimental Study
In this section, we ask two questions. First, what is the impact on adherence of leveraging the methods of this paper to optimize inspection policies? Second, what is the cost of the inspection service constraints in terms of adherence? We describe the compared policies in Sec. 6.1 and study the impact of different planning policies on reward and computation time, both in synthetic (Sec. 6.2) and real data from CDPH (Sec. 6.3) domains.
@@ 6.1 Planning Policies
  • (Rdm, Opt, $\boldsymbol{\leq}$ 1): The action windows are distributed at random, each arm needs to be pulled at most once per period, and the method of Sec 4.2 is used to optimize rewards.
  • (Opt, Opt, $\boldsymbol{\leq}$ 1): The action windows are optimized and each arm needs to be pulled at most once per period. We optimize windows, and then we model the RMAB as Sec 4.2.
@@ 6.2 Synthetic Domain
6.2.1 Data Preparation and Setup
@@ 6.3 Food Establishment Inspection Domain
Using inspection data from the Chicago Data Portal Chicago Data Portal (2023), we implement a realistic RMAB setting.
6.3.1 Data and Setup
Since 2010, CDPH has published every food establishment inspection result on the Chicago Data Portal Chicago Data Portal (2023). The Food Inspection Dataset is a tabular dataset with 17 attributes for each establishment including license number, address, etc. The inspection results are shown in the “Violations” column: 0 means no violations and pass, 1 means violations appear and 2 means pass with conditions. In the experiment, both 0 and 2 are merged into a single good state and 1 is the bad state.
Inferring Transitions
To create a realistic instance, we must infer the transition probabilities from the inspection trajectories for each arm. Because of the small number of inspections per establishment, it is impractical to estimate a transition model for each food establishment. Thus, we used all data to train one single model to learn the transitions for all establishments. A neural network with 2 MLPs is trained to predict the transition matrix $P^{(0)}_{i}$ to maximize the log-likelihood of the data given the transitions. The detailed architecture of the neural network is provided in the supplement.
Data Preparation
The inputs to the model are a series of features of establishments. We use the same features as previous work on predicting food establishment risks  Schenk Jr. et al. (2015), which combines data of business licenses, food inspections, crime, garbage complaints, sanitation complaints, weather and sanitation information.
Loss
@@ 6.3.1 Data and Setup
The fit to real-world data is excellent compared with the XGBoost model with an AUC of 0.67 in the Schenk Jr. et al. (2015).

### v2

# Optimizing Urban Service Allocation with Time-Constrained Restless Bandits

@@ Abstract
Municipal inspections are an important part of maintaining the quality of goods and services. In this paper, we approach the problem of intelligently scheduling service inspections to maximize their impact, using the case of food establishment inspections in Chicago as a case study. The Chicago Department of Public Health (CDPH) inspects thousands of establishments each year, with a substantial fail rate (over 3,000 failed inspection reports in 2023). To balance the objectives of ensuring adherence to guidelines, minimizing disruption to establishments, and minimizing inspection costs, CDPH assigns each establishment an inspection window every year and guarantees that they will be inspected exactly once during that window. Meanwhile, CDPH also promises surprise public health inspections for unexpected food safety emergencies or complaints. These constraints create a challenge for a restless multi-armed bandit (RMAB) approach, for which there are no existing methods. We develop an extension to Whittle index-based systems for RMABs that can guarantee action window constraints and frequencies, and furthermore can be leveraged to optimize action window assignments themselves. Briefly, we combine MDP reformulation and integer programming-based lookahead to maximize the impact of inspections subject to constraints. A neural network-based supervised learning model is developed to model state transitions of real Chicago establishments using public CDPH inspection records, which demonstrates 10% AUC improvements compared with directly predicting establishments’ failures. Our experiments not only show up to 24% (in simulation) or 33% (on real data) objective improvements resulting from our approach and robustness to surprise inspections, but also give insight into the impact of scheduling constraints.
@@ 1 Introduction
Restless multi-armed bandits (RMABs) (whittle_1988) are almost a natural fit for this problem. They describe a sequential decision problem where an agent (the city) acts on a large population of independently evolving Markov decision processes (the establishments), which describe each establishment’s propensity to stay in the inspection-passing state. While RMABs are highly suitable for maximizing the inspection-passing probability for each establishment, existing approaches (e.g.,  (deadline; fair_bandit; li2023avoiding)) fail to support the constraint structure required for establishment scheduling, where each arm’s pulls must satisfy an ex-post frequency constraint.
@@
  • Empirical evaluation and impact/robustness assessment. For the CDPH, we develop a machine-learning approach to estimate the parameters of each establishment based on historical data. We evaluate reward and computation time in both real and synthetic data. In our tested cases, we find that both explicitly modeled constraints and optimized action windows bring substantial advantages compared with ad hoc methods for handling constraints. Considering some urban emergent service would break the original schedulings, we evaluate our method’s robustness by adding unexpected mandatory actions and find our method is more robust than ad hoc plannings.
  • Machine-learning model predicting transitions of MDP. There is one previous work predicting the food inspection failures directly using machine-learning models (foodforecast). For the purpose of building MDPs for our RMAB, we instead train a neural network to predict the state transition probabilities. We found the model not only aligns with the requirements of RMAB, but also has the ability to predicting the inspection failures. It improves the AUC number by around 10% compared with the XGBoost model proposed by foodforecast.
@@ 2 Related Work
In 2015, CDPH leveraged historical food inspection data and trained a supervised learning model to predict the probability that an inspection would uncover a critical violation (foodforecast).  kannan2019hindsight independently analyzed the impact of prediction-driven scheduling. However, such models only consider one-shot predictions for critical violations and do not include the sequential aspect of scheduling. Fairness is of substantial interest in the provision of municipal services  (singh2022fair). We consider fairness outside the scope of this paper, but a potentially interesting direction for future work.
Restless multi-armed bandits (RMABs). RMABs are PSPACE-hard in the worst case, but whittle_1988 showed that a subclass of them, so-called indexable RMABs, admit an efficient asymptotically optimal solution. Particularly relevant classes of indexable RMABs are those that extend the machine maintenance problem families (some_family), and scheduling problems for sensors (aoi), wireless transmission (wireless_trans), and health interventions (mate2020collapsing). These RMABs are structured so the state of each process declines if it is not acted on, and differ in the details of action effect and what information is observed with or without an action. This work aims to develop techniques to integrate action constraints into RMABs of these types.
RMABs with Constraints. Several RMAB models have included constraints. In a project applying RMABs to assist maternal and child health via phone calls, a “sleeping period” for arms was enforced after they were pulled by the Whittle index heuristic (field_study) (See 4.2). It appears to have been enforced in an ad hoc manner, by blocking pulls that would have violated the constraint.  deadline deployed RMABs in a deadline scheduling setting and integrated the deadline constraints by adding dummy arms. Fairness is another setting where constraints can arise.  fair_bandit introduced the ProbFair policy, ensuring a strictly positive lower bound on the probability of being pulled at each time step while still satisfying the budget constraints.  li2023avoiding’s SoftFair balances the trade-off between the goals of having resources uniformly distributed and maximizing cumulative rewards. They guarantee a long-term probability of each arm being pulled whereas ours ensures pulling frequencies for each arm strictly in a period. To the best of our knowledge, we are the first to consider action window and frequency constraints strictly for recurring tasks in a period.
@@ 3 Preliminaries
An RMAB consists of $N$ binary action MDPs (arms). We define the $i$th two-action MDP (10.5555/528623) as a tuple $(\mathcal{S}_{i},\mathcal{A},P_{i},R_{i},s^{(0)}_{i},\gamma)$. The discount factor $\gamma$ and action space $\mathcal{A}=\{0,1\}$ are fixed across all MDPs. When the action 1 (resp. 0) is taken on an arm at time $t$, we refer to that arm as active (resp. passive). The rest are arm specific: $\mathcal{S}_{i}$ is the state space, $P_{i}:\mathcal{S}_{i}\times\mathcal{A}\rightarrow\Delta\mathcal{S}_{i}$ is the transition function, $s^{(0)}_{i}$ is the start state, and $R_{i}:\mathcal{S}_{i}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. Because there are only two actions, the transition function $P_{i}$ can be decomposed into an an active transition $P^{(1)}_{i}:\mathcal{S}_{i}\rightarrow\Delta\mathcal{S}_{i}$ and a passive transition $P^{(0)}_{i}:\mathcal{S}_{i}\rightarrow\Delta\mathcal{S}_{i}$.
@@ 3.1 Whittle Indices
General RMABs have an exponentially large state space and a combinatorially large action space. The Whittle index method provides tractability for some classes of RMABs (whittle_1988). It works by computing a “benefit of acting” for each arm, called the Whittle index. The Whittle index heuristic then acts on the $k$ arms with highest Whittle indices.
@@
The goal is to identify the smallest subsidy $m$ such that, for the current state $s_{i,t}$, the long-term rewards for the passive and active actions are the same. To formalize this, we first define the $Q$ function for arm $i$ under subsidy $m$:
@@
Definition
The Whittle index for state $s_{i,t}$ is the smallest m which makes it equally optimal to take the active and passive actions:
@@
For the Whittle index heuristic to have asymptotic optimality guarantees, each arm must satisfy a technical condition called indexability (whittle_1988). Intuitively, indexability says that, as $m$ increases, the optimal action can only switch to passive and cannot switch back to active. Let $W_{i}^{(m)}$ be the set of states for which $Q_{i}^{(m)}(s_{i,t},a=0)\geq Q_{i}^{(m)}(s_{i,t},a=1)$, i.e., the passive action has an equal or higher return than the active action.
Definition
(Indexability). An arm is said to be indexable if $W_{i}^{(m)}$ is non-decreasing in m, i.e., for any $m_{1},m_{2}\in\mathbb{R}$ such that $m_{1}\leq m_{2}$, we have $W_{i}^{(m_{1})}\subseteq W_{i}^{(m_{2})}$. An RMAB is indexable if every arm is indexable.
@@ 4.1 Motivating Inspection RMAB
Motivated by the food establishment setting, we define a model RMAB to which we will add action constraints. This RMAB can be viewed as a collapsing bandit (mate2020collapsing) or a resetting bandit (resource), and both have indexability guarantees. Each establishment has an unobserved binary state that is either 1 (i.e., inspection passing) or 0 (i.e., inspection failing). When we act on the establishment, we assume that it is restored to the passing state and define the reward function to be 1 for each time period the establishment is in the passing state and 0 otherwise. We consider time periods as months—each establishment needs to be inspected once a year and will have a two-month period where this inspection can occur.
As the true states are not directly observable, each arm is a partially observed Markov decision process (POMDP) (pomdp). We can rewrite the POMDP as a fully observed belief-state MDP, allowing for direct representation as an RMAB.
@@ 4.2 Action Windows and MDP Encoding
We use action windows as an exemplar for the family constraints where the constraint can be directly encoded into the RMAB structure, i.e., a vanilla RMAB with an action window constraint can be rewritten as a vanilla unconstrained RMAB a different arm structure. This is in some sense the ideal way to add constraints—we can apply whatever existing state-of-the-art algorithm directly.
@@
As shown in Figure 1, such an encoding increases the number of states in the MDP. The number of states is increased by a factor of $O(LM)$, where $L$ is the number of counter values required to track when the window is active, and $M$ is the total number of actions allowed during the window. For the motivating RMAB, this increase is by a factor of 14. But such encoding is also applicable to more complex situations: An arm has multiple action windows and multiple inspections allowed per window. By design, the Whittle heuristic ensures no arms outside the action window are selected and no ill effect even if selected by accident. We provide a precise description of the new MDP in the appendix. All new transitions are deterministic (probability = 1).
@@ 4.3 Frequency Constraints and Lookahead
Maximizing the sum of Whittle indices without additional frequency constraints OR with the constraint that each arm must be pulled exactly once during the lookahead window can be reduced to a weighted $b$-matching (doi:b_matching).
@@
In practice, it is convenient to solve this lookahead problem as an IP. We can do so with $NL$ binary variables $a_{i,t}$ (where $L$ is the length of the action window) and $T+N$ constraints. Because the polynomial tractability of weighted $b$-matching arises from total unimodularity (b-matching), the IP can be solved very quickly via its LP relaxation. However, polynomial tractability is lost when sufficiently complex minimum and maximum number of pulls are added as additional constraints as the problem becomes equivalent to weighted bipartite $b$-matching (wbm).
@@ 5 Action Window Optimization
How do we assign windows according to the virtual inspection sequence? Let $a_{i,t}$ be the encoding of the virtual inspection sequence, where $a_{i,t}=1$ if arm $i$ is pulled at time $t$. A naive way is to assign arm $i$ with window [t,t+W1][t,t+W-1] if $a_{i,t}=1$. However, such assignments are easily predictable and establishments would be able to prepare effectively. We need to design a way in which establishments get inspected with a probability of $1/W$ on each day in the window.
We can do so with a linear program (LP). We define variables $f_{t,t^{\prime}}$, which indicate the proportion of arms with virtual inspection time $t$ that are put in inspection window $t^{\prime},t^{\prime}+W-1]$. We denote the number of arms with assigned inspection $t$ that are put in the window starting at $t^{\prime}$ as $g_{t,t^{\prime}}$. Thus, $g_{t,t^{\prime}}=\sum\limits_{i=1}^{N}a_{i,t}f_{t,t^{\prime}}$. For the objective, we use
@@
    The virtual action assignment must occur during the window: $f_{t,t^{\prime}}\neq 0$ if and only if $0\leq t-t^{\prime}\leq W-1$.
@@
Using the $f_{t,t^{\prime}}$ output of the LP, we sample windows using any categorical sampling procedure. At this stage, the virtual inspection can be discarded and individual inspections planned as if the windows were given (i.e., not even the system operator knows when the virtual inspections were planned to occur). In practice, unplanned inspections (i.e., due to complaints) will happen constantly, meaning replanning is necessary to satisfy constraints.
@@ 6 Experimental Study
In this section, we ask three questions. First, what is the impact on adherence of leveraging the methods of this paper to optimize inspection policies? Second, what is the cost of the inspection service constraints in terms of adherence? Third, how do different methods degrade when unplanned complaint inspections are occuring? We describe the compared policies in Sec. 6.1 and study the impact of different planning policies on reward and computation time, both in synthetic (Sec. 6.2) and real data from CDPH (Sec. 6.3) domains.
@@ 6.1 Planning Policies
[…]
@@ 6.2 Synthetic Domain
Data Preparation and Setup
@@ 6.3 Food Establishment Inspection Domain
Using inspection data from the Chicago Data Portal (foodinspectiondata), we implement a realistic RMAB setting.
Data and Setup
Since 2010, CDPH has published every food establishment inspection result on the Chicago Data Portal (foodinspectiondata). The Food Inspection Dataset is a tabular dataset with 17 attributes for each establishment including license number, address, inspection results, etc. A neural network with 2 MLPs is trained to predict the transition probabilities. More details of data and model structures are shown in the appendix.
Loss of Model Training
@@
The fit to real-world data is excellent compared with the XGBoost model with an AUC of 0.67 in the (foodforecast).